40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 99 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  12094   Wed Apr 27 15:04:47 2016 SteveUpdateendtable upgradeCleaning ETMX vacuum dirty window

It looks very promising.

 

Attachment 1: 1cETMX-Tcmp.jpg
1cETMX-Tcmp.jpg
  12093   Wed Apr 27 14:06:31 2016 ranaSummaryGeneralmeeting notes
  1. Gautam will get help from Johannes and finish EX table by Monday.
  2. Steve will spend a day this week with Johannes on Green Monster bakeout.
  3. Q to analyzed green PDH servo and design demod low pass. Should we use the double LC notches to notch the 2f product? What's the demod filter attenuation requirement?
  4. Koji will make a drawing of the ruby suspension standoff prism and post into the elog so that Steve can get some quotes next week.
  5. Rana to implement 40m configuration in FOGprime17 and analyze RoC matching of ETMs. Get Antonio's help to analyzed SRC stability. Maybe use PyKat and Finesse since Antonio knows that stuff.
  6. Give OCXO boxes to WB refcav people. Rana get Rich to make another couple of boxes for 40m PMC, FSS, IMC.
  7. Rana/Koji get EKG to make specs and procure some new folding mirrors for the PRC/SRC. Make them a bit concave and dichroic.
  12092   Wed Apr 27 09:45:56 2016 ranaUpdateSUSspare SOS tower

Bah, we need ruby slippers for all future suspensions. Prism with curved backside and smooth grooves.

No aluminum, no cry.

Quote:

Earth quake stops need viton tips.

Wirestandoffs are still aluminum.

 

  12091   Wed Apr 27 09:05:10 2016 SteveUpdateGeneralAP viewport

                   Sad situation

    The anti-symmetric port

spider webs fly in the wind

Attachment 1: APg.jpg
APg.jpg
Attachment 2: AP.jpg
AP.jpg
Attachment 3: APspiderWebs.jpg
APspiderWebs.jpg
  12090   Tue Apr 26 23:19:42 2016 gautamUpdateendtable upgradeGreen aligned to arm - high order mode flashes seen

Attachment #1

Layout as of today. Most of the green path is done. The Green REFL PD + PZT mirrors have not been hooked up to their respective power sources yet (I wonder if it's okay to start laying cables through the feedthroughs on either end of the table already, or if we want to put whatever it is that makes it airtight eventually in first?). A rough power budget has been included (with no harmonic separator just before the window), though some optimization can be done once the table is completely repopulated.

Attachment #2

A zoomed-in version of the REFL path.

Some general notes:

  1. I've tried to use the custom 3/4" O.D. posts + baseplate arrangement wherever possible (only 1 steering mirror is on a 1" post clamped with a fork to the table because of space constraints). Where the baseplates could not be bolted onto the table directly, I've used Newport SS Dogs to do the job.
  2. I checked for continuity between the PZT outer case and the table top with a multimeter, and found none. So I chose to leave the Thorlabs baseplates in place. For the REFL PD, I've used an insulating baseplate given to me by Steve.
  3. I've used some custom length 3/4" O.D. posts to get the beam up to the right height (~4.75") just before sending the green beam in. The beam height is 4" elsewhere.
  4. I was playing around with positioning the harmonic separator immediately before the vacuum chamber window - I found that there is a substantial amount of green light that is reflected, though there doesn't seem to be any IR leaking through. The mirror was labelled Y1-1037-45P, which is a code for CVI mirrors, though I believe it is a LaserOptik product and that we have a couple of other such mirrors in the optics cabinet - though they are all 1". This document suggests that from the back side, there should be <0.1% reflection of green while on the front side it should be < 3%. I will have to hunt a little more for the specs, and measure the powers to see if they match the previously quoted numbers. In any case, I'll have to think of how to separate the (unwanted) reflected green and the transmitted IR from the cavity in the IR transmon path.
  5. There are some minor changes to the planned layout posted here - I will update these in due course once the Transmon path and Oplev have been set up.

I am closing the PSL shutter and the EX laser shutters for the night as I have applied a layer of first contact to the window for cleaning purposes, and we don't want any laser light incident on it. It may be that the window is so dirty that we may need multiple F.C. cleaning rounds, we will see how the window looks tomorrow...

 

Attachment 1: IMG_2219.JPG
IMG_2219.JPG
Attachment 2: IMG_2220.JPG
IMG_2220.JPG
  12089   Tue Apr 26 15:22:35 2016 SteveUpdateendtable upgradeCleaning ETMX vacuum dirty window

Gautom is progressing with the layout nicely. The X-arm transmission window have not seen cleaning for decades. This should be the time to do it. Here is picture of dirtiness.

It is not that simple... How much effort should we put in it? The hole table with 1W inno laser plus... set up now about ~500 lbs  We can pull it off carefully, but it is not risk free.

We should look at our other signal port windows! Gautom's long reach able him to do the first contact cleaning without moving anything. It is great!
 

Attachment 1: ETMX-Tvp.jpg
ETMX-Tvp.jpg
Attachment 2: ETMX-TvpDetail.jpg
ETMX-TvpDetail.jpg
  12088   Mon Apr 25 11:07:06 2016 SteveUpdateSUSspare SOS tower

Earth quake stops need viton tips.

Wirestandoffs are still aluminum.

Attachment 1: ETMXreplacment.jpg
ETMXreplacment.jpg
  12087   Fri Apr 22 13:58:13 2016 SteveUpdatePEMleaky roof is fixed

Dan sealed the leak today.

 

Attachment 1: leakyRoof_(2).jpg
leakyRoof_(2).jpg
  12086   Thu Apr 21 15:12:38 2016 SteveUpdateVACRGA is not working

 

Quote:

Steve pointed out that in the aftermath of the Nitrogen running out a couple of times last week, the RGA had shut itself off thinking that there was a leak and so it was not performing the scheduled scans once a day. So the data files from the scheduled scans were empty in the /opt/rtcds/caltech/c1/scripts/RGA/logs directory. The wiki page for getting it up and running again is up-to-date, but the script RGAset.py did not exist on the c0rga machine, which the RGA is communicating with via serial port. I copied over the script RGAset.py from rossa to c0rga and ran the script on that machine - but the error flags it returned were not all 0 (indicating some error according to the manual) - so I edited the script to send just the initialize command ('IN0') and commented out the other commands, after which I got error flags which were all 0. After this, I ran a manual scan using 'RGAlogger.py', and it appears that the RGA is now able to take scans again - I'm attaching a plot of the scan results. We've saved this scan as a reference to compare against after a few days. 

Our last RGA scan is from February 14, 2016  We had a power outage on the 15th

Gautom has not succeded  reseting it. The old c0rga computer looks dead. Q may resurrect it, if he can?

  12085   Thu Apr 21 14:25:52 2016 gautamUpdateendtable upgradeGreen light recovered

I've made progress on the new layout up to the doubling oven. After doing the coarse alignment with the diode current to the NPRO at ~1A, I turned it back up to the nominal 2A. I then rotated the HWP before the IR Faraday such that only ~470mW of IR power is going into the doubler (the rest is being dumped on razor beam dumps). After tuning the alignment of the IR into the doubling oven using the steering mirror + 4 axis translation stage on which the doubling oven is mounted, I get ~3.2mW of green after the harmonic separator and a HR mirror for green. The mode looks pretty good to the eye (see attachment #1), and the conversion efficiency is ~1.45%/W - which is somewhat less than the expected 2%/W but in the ballpark. It may be that some fine tweaking of the alignment + polarization while monitoring the green power can improve the situation a little bit (I think it may go up to ~4mW, which would be pretty close to 2%/W conversion efficiency). The harmonic separator also seems to be reflecting quite a bit of green light along with IR (see attachment #2) - so I'm not sure how much of a correction that introduces to the conversion efficiency. 

While doing the alignment, I noticed that some amount of IR light is actually transmitted through the HR mirrors. With ~500mW of incident light at ~45 degrees, this transmitted light amounts to ~2mW. Turns out that this is also polarization dependant (see attachment #3) - for S polarized light, as at the first two steering mirrors after the NPRO, there is no transmitted light, while for P-polarized light, which is what we want for the doubling crystal, the amount transmitted is ~0.5%. The point is, I think the measured levels are consistent with the CVI datasheet. We just have to take care find all these stray beams and dump them.

I will try and optimize the amount of green power we can get out of the doubler a little more (but anyway 3mW should still be plenty for ALS). Once I'm happy with that, I will proceed with laying out the optics for mode-matching the green to the arm.

Attachment 1: IMG_6567.JPG
IMG_6567.JPG
Attachment 2: IMG_6568.JPG
IMG_6568.JPG
Attachment 3: CVI_reflectivity.jpeg
CVI_reflectivity.jpeg
  12084   Wed Apr 20 14:36:45 2016 SteveUpdatePEMdusty lab

It is worth wiping table top covers. Use isopropanol soaked lint free wipes.

 

Attachment 1: AP_accm.jpg
AP_accm.jpg
Attachment 2: wiped_AP.jpg
wiped_AP.jpg
Attachment 3: temp-partic_6months.png
temp-partic_6months.png
Attachment 4: 3_years_PEM.png
3_years_PEM.png
  12083   Tue Apr 19 18:37:29 2016 gautamUpdateendtable upgradeLaser swap + optical layout

Summary of work done over the last two days

  1. Lightwave NPRO + controller moved to PSL table
    • ​​The interlock is not connected to the controller
    • Controller is not powered
  2. Innolight NPRO + controller installed at endtable
    • ​​​​Interlock has been connected
    • For initial alignment purposes, I'm running it at an injection current of 1.000A (~50mW of IR out of the NPRO)
    • Temperature of crystal set to 31.66 degrees in anticipation of operation in the nominal state
  3. Laying out optics
    • ​​Given that the mode out of the NPRO is different from that from the Lightwave, the mode-matching had to be re-done
    • Attachment #1 shows the mode-matching solution being implemented
    • Current state - I've placed all the optics up to and including the doubling crystal + oven. Alignment through IR Faraday is pretty good, QWP+HWP angles optimized to maximize transmission through the Faraday (<10% loss). Oven has been hooked up to temperature controller, and is currently set to 36.3 degrees. Coarse alignment into doubling crystal done at lower power. Even with the low IR power, I am able to see some green. It remains to turn the injection current up and do the fine alignment + lens position tweaking to maximize the green power from the doubling crystal - with ~1W of power, assuming 2%/W SHG efficiency, we should be seeing 20 mW of green (which is probably way too much)

Immediate next steps:

  1. Some optimization to be done with regards to beam dumps for rejected beam from IR Faraday. Also double check to make sure that the reflected beam from L1 doesn't go back directly to the laser (at the moment it doesn't, is there a standard way to do this? I was trying to have the lens as close to normal incidence as possible, but I may not have been entirely successful which is why the reflected beam is not going straight back at the moment).
  2. Optimize mode-matching into the doubling crystal
  3. Once the desired green mode is obtained, continue with the rest of the layout
  4. Update CAD drawing to reflect new layout

 

Attachment 1: IR_modematch_19April2016_2.pdf
IR_modematch_19April2016_2.pdf
  12082   Tue Apr 19 10:58:53 2016 SteveUpdatesafetysafety scans at PSL & tables

The PSL had one 1064 nm beam to be blocked around the north east side. The end enclosures are fine.

 

 

  12081   Mon Apr 18 00:29:00 2016 gautamUpdateGeneralBeam profiling + injection current scan

Summary

I've finished up the remaining characterization of the repaired 1W Innolight NPRO - the beamscan yielded results that are consistent with an earlier beam-profiling and also the numbers in the datasheet. The output power vs diode current plot is mainly for diagnostic purposes in the future - so the plot itself doesn't signify anything, but I'm uploading the data here for future reference. The methodology and analysis framework for the beamscan is the same as was used here.

Attachment #1 - Beam-scan results for X-direction

Attachment #2 - Beam-scan results for Y-direction

Attachment #3 - Beam profile using fitted beam radii

Attachment #4 - Beam-scan data

Attachment #5 - Output power vs Injection current plot

Even though I remember operating at a diode current of 2.1A at some point in the past, while doing this scan, attempting to increase the current above 2.07A resulted in the "Clamp" LED on the front turning on. According to the manual, this means that the internal current limiting circuitry has kicked in. But I don't think this is a problem as we don't really even need 1W of output power. This is probably an indicator of the health of the diode as well?

Attachment #6 - Output power vs Injection current data

It remains to redo the mode-matching into the doubling oven and make slight modifications to the layout to accommodate the new laser + beam profile. 

I plan to do these in the morning tomorrow, and unless there are any objections, I will begin installing the repaired 1W Innolight Mephisto on the X endtable tomorrow (18 April 2016) afternoon. 

 
Attachment 1: BeamScan_x.pdf
BeamScan_x.pdf
Attachment 2: BeamScan_y.pdf
BeamScan_y.pdf
Attachment 3: ZScan.pdf
ZScan.pdf
Attachment 4: BeamScan.mat
Attachment 5: Innolight_Current_Scan.pdf
Innolight_Current_Scan.pdf
Attachment 6: Innolight_Current_Scan.mat
  12080   Fri Apr 15 23:11:49 2016 gautamUpdateGeneralInnolight 1W moved to SP table

I have moved the 1W Innolight + controller from the PSL table to the SP table for beam profiling.

  12079   Fri Apr 15 18:38:12 2016 gautamUpdateendtable upgradeLightwave health check - NO IMPROVEMENT

I re-measured the power levels today.

We have ~205mW out of the NPRO, and ~190mW after the Faraday. It doesn't look like the situation is going to improve dramatically. I'm going to work on a revised layout with the Innolight as soon as I've profiled the beam from it, and hopefully, by Monday, we can decide that we are going ahead with using the Innolight.

  12078   Fri Apr 15 18:35:57 2016 gautamUpdateGeneralNew AUX laser measurements

I've performed the temperature sweep of PSL vs Innolight 1W AUX laser.

  • I followed the procedure in this elog - started by turning of FSS and FSS Slow servos, closed the PSL shutter, noted down the value of PSL temperature
  • As noted in elog 3759, there are multiple temperatures at which a beat can be found. I recorded all that I could find. The IR beat frequency was < 20MHz at the temperatures recorded (and had an amplitude of a few dBm, but I used a 20dB coupler to look at the signal on the HP spectrum analyzer
  • The PMC unlocked each time I changed the PSL temperature, but the PMC autolocker worked for me every time
  • We should use curve 3 in attachment 1, it is the most reliable set of temperatures at which a beat can be found
  • PSL diode current was 2.100A, AUX laser diode current was 2.001A
  • Attachment 2 is the data

It remains to measure the output power vs diode current, and the beam profile. I will do the latter on the SP table where there is a little more space. Because we have 1W from this NPRO, the knife-edge method requires a power meter that has a large dynamic range and is sensitive enough to profile the beam accurately. After consulting the datasheets of the power meters we have available (Scientech, Ophir and Coherent) together with Koji, I have concluded that the Coherent calorimeter will be suitable. Its datasheet claims it can accurately measure incident powers of up to 100uW, although I think the threshold is more like 5-10mW, but this should still be plenty to get sufficient resolution for a Gaussian intensity profile with peak intensity of 1W. We also checked that the maximum likely power density we are likely to have during the waist measurement process (1W in a beam of diameter 160um) is within the 6kW/cm^2 quoted on the datasheet.

Attachment 1: PSL_AUX_TEMP_SCAN.pdf
PSL_AUX_TEMP_SCAN.pdf
Attachment 2: PSL_AUX_TEMP_SCAN.mat
  12077   Fri Apr 15 03:02:44 2016 ericqUpdateGeneralNew AUX laser measurements

The free running PSL+AUX beat frequency noise spectrum has been measured via PLL. AUX laser PZT PM and AM responses were measured too. 


Rough notes about these measurements:

Laser -> QWP -> HWP -> PBS -> 10% BS -> Beat
3.4Vpp out of PD, (40% contrast)
20dB Coupler, output to analyzer, coupled output to Mixer (-a few dBm, didn't check specifically)
Mixer: ZP-3+, BLP-5.1 at output
LO: OCXO @ 36MHz 13dBm->5dB Att-> +8dBm LO at Mixer

Got ~65mVpp out of Mixer

Mixer out -> SR560, LP 3Hz, G=500 -> Pomona Summing node -> Laser PZT
~30kHz UGF ~30 deg phase

Spectra, OLG via SR785 taken with free running PSL, anthropomorphic temperature servo. Data sheet calibration used for PZT. SR560 output noise dominates over analyzer, mixer, PD. Spectrum looks ok, I think.

PM measured with AG4395. High impedance probe used for laser PZT, otherwise couldn't lock. PM calibrated via mixer voltage span for fringe-to-fringe. 

PSL beam blocked, AUX power increased to read 8.0V, AM measured with AG4395.

AM/PM doesn't look to dissimilar to old measurements on wiki. ~230kHz looks like a fine modulation freq. 

Still to be done to AUX laser:
- joint PSL/AUX temperature sweeps
- Output power vs. diode current
- Beam profile

Attachment 1: aux_innolight_freqPSD.pdf
aux_innolight_freqPSD.pdf
Attachment 2: aux_innolight_AMPM.pdf
aux_innolight_AMPM.pdf
Attachment 3: aux_innolight_data.zip
  12076   Thu Apr 14 17:30:18 2016 ericqUpdateGeneralPLL measurement ongoing

Just a heads up that some equipment is hooked up at the PSL table for the repaired AUX laser PLL measurement, I plan to continue with it tonight.

I've taken a few spectra that, along with the PZT coefficient from the repair sheet, that suggest the noise level is ok (incoherent sum of AUX and PSL at about ~3e4 / f Hz/rtHz), but calibrated plots, etc. will follow in time.

  12075   Wed Apr 13 18:25:07 2016 gautamUpdateendtable upgradeLightwave health check

[Koji,gautam]

Lightwave NPRO information:

Model: 126-1064-700

Serial Number: 337

Manufactured: December 1998!!

Details of checks performed:

Koji tuned the parameters on the laser controller and we observed the following:

  1. Turning "ADJ" to +10 and the pumping current all the way up to the maximum (2.62A) allowed us to recover an output power of 300mW, at a laser crystal temperature of ~45degrees
  2. The output power increased almost monotonically as a function of the laser crystal temperature - why? We were able to see powers as high as 250mW (at ADJ=0) for the maximum crystal temperature of ~60 degrees.
  3. We checked that we could believe the readout of the power meter by measuring the power using the Scientech power meter - we saw ~270mW after the Faraday with this meter, accounting for ~10% loss through the Faraday, this corresponds to an output power of 300mW (all this was done at ADJ=+10, DC=2.62A). I suspect that the display is dodgy though, because changing the Diode Current from 2.52A to 2.62A increased the output power by almost 100mW, which seems hard to believe?
  4. The Lightwave NPRO does not have heat dissipation fins attached - could this be affecting the power output somehow? In any case, this has to be rectified. So if we decide to keep the Lightwave NPRO, the layout will still need minor changes to accommodate the heat fins. Steve, do we have these in hand?

Way forward

Ericq has begun the characterization of the repaired Innolight. We checked that it outputs 1W of power. We will now have to perform the following measurements:

  • Frequency noise using PLL
  • AM/PM response of the PZT
  • Laser power output as a function of diode current - this will be useful for diagnostic purposes in the future
  • AUX temperature vs PSL temperature at which beatnotes can be found
  • Waist measurement - the mode matching and optical layout upstream of the doubling oven at least will have to be modified significantly

All of these will have to be done before installing this laser at the endtable.

I believe the consensus as of now is to go ahead with carrying out the above measurements. Meanwhile, we will keep the Lightwave NPRO on and see if there is some miraculous improvement. So the decision as to whether to use the Innolight is deferred for a day or two.

  12074   Wed Apr 13 11:00:28 2016 SteveUpdateendtable upgradeX endtable repopulation

ETMX optical table is grounded to ETMX chamber through 1 Mohms

The doubling oven temp controller is installed to reach its cable.

 

  12073   Wed Apr 13 00:56:07 2016 gautamUpdateendtable upgradeX endtable repopulation

Over the last couple of days, I've been working on restoring the optical layout on the X-endtable. Some notes about the status as of today:

Lightwave NPRO output power

The output power from the lightwave NPRO is about 210mW (as measured with the calorimeter). This is significantly lower than the value of ~300mW reported in this elog. It may be that the laser crystal temperature has changed compared to that measurement, but the "ADJ" parameter is at 0, both today and in that measurement. The laser has also been on for more than a day now, that should be sufficient time for the crystal to equilibriate to its final operating state? Is such a large change in output power possible just because of a change in laser crystal temperature? Or did the laser really lose ~1/3rd of its output power over the last two months?

Alignment into IR Faraday, and changes to the planned layout

I've set up the layout until steering the beam through the IR faraday. The input power into the IR Faraday is ~210mW. The output power is ~186mW, after optimizing the angle of the HWP. These numbers seem consistent with what I had reported in this elog (although this was for the Innolight NPRO). The alignment looks reasonably good to the eye as well.

I've made one change to the planned layout (latest version here). Y1 is now a 2" 99% reflective for S polarization beam splitter, instead of a 1" HR mirror. I made this change because we want some light from the NPRO to be transmitted through this optic to couple into the fiber eventually, for the IR beat. I measured the transmitted power to be ~1.5mW, which is around what we were coupling into the fiber before, and should suffice now. The Lightwave NPRO datasheet (page 4) suggests that the polarization of the output of the laser is S, and the measured power before and after this optic suggests that it is working as advertised. This means that HWP 1 also has to be moved downstream (to rotate the polarization so as to maximize transmission through the IR faraday). Space constraints meant that I could not mount HWP 1 on the baseplate+3/4" OD post assembly which is what we want where possible on the new table, so for this optic, I used a 1" OD post and a fork. There may be a couple of other optics in the final layout where space constraints dictate we compromise in this way.

I've also installed beam dumps for the rejected light from the Faraday. For now, these are the old beam dumps. They looked reasonably intact. I believe we have a bunch of new beam dumps on hand as well, so these can be swapped out if deemed necessary.

Cleaning of optics

All the optics are being cleaned using first contact before being installed on the table. 

As I found out the hard way, it is not a good idea to clean small optics like half-wave plates while in their mounts. The first contact tends to bond to the frame while drying, and doesn't come off cleanly. Koji helped me clean the offending pieces (he used tweezers to manually remove the residual first contact, and then some acetone to clean up any remaining residue). Subsequently, he re-cleaned these optics, again using first contact, but this time being careful not to extend all the way out to the edge of the optic. The idea is to cover as much area as possible with first contact, while staying clear of the edge. This approach worked reasonably well.


The next major step is to achieve optimal alignment into the doubler. I've placed the doubler on the table in it's approximate final position, I wanted to make sure the enclosure support wasn't in the way (it isn't). The cable from the oven won't run all the way to the Thorlabs temperature controller in it's usual place, we need to either extend the cable, or figure out a new place where we can keep the temperature controller.

  12072   Tue Apr 12 22:41:00 2016 KojiUpdateIOOPMC/IMC aligned, WFS offset adjusted

Did it again.

PMC Trans ~0.739
IMC Trans ~15000

  12071   Tue Apr 12 09:14:57 2016 SteveUpdateSUSRuby wire - v - groove cut pictures

The ruby wire standoff V groove cuts are looking good.

I will request free sample of  sapphire prizm where one side would have SOS's R cylindrical surface.

The present plan to have the v-groove on this prism.

 

Attachment 1: Sapphire_prism_wire_standoff.JPG
Sapphire_prism_wire_standoff.JPG
  12070   Mon Apr 11 17:03:41 2016 SteveUpdateCalibration-Repair1W Innolight repair completed

The laser is back. Test report is in the 40m wiki as New Pump Diode Mephisto 1000

It will go on the PSL table.

  12069   Mon Apr 11 16:06:30 2016 ericqUpdateLSCDRFPMI Data Archived

I have copied over the complete frame files from two DRFPMI lock acquisitions + locks to /frames/archive. The data should be safe from the wiper script here.

One, under the subfolder DRFPMI_Mar29_cal is the lock where the CAL-DARM channel is properly calibrated at GPS time 1143274087.

The other lock, under DRFPMI_MAR29_nocal, does not have the calibration set up yet, but was a much quicker acquistion (<2 min from ALS acquisition to DRFPMI) and longer lock (~8min).

  12068   Fri Apr 8 09:13:20 2016 SteveUpdateendtable upgradeETMX-T beam height

X arm resonating after alignment, beam height on ETMX optical table ~4.75"

 

Attachment 1: ETMXalignedT.png
ETMXalignedT.png
Attachment 2: ETMXF.jpg
ETMXF.jpg
  12067   Thu Apr 7 14:30:58 2016 SteveUpdateendtable upgradeETMX table height

 

Quote:

Steve has finished installing the enclosure on the new endtable. So Eric and I decided to try and lock the X arm and measure the beam height of the transmitted IR beam relative to the endtable. We initially thought of using POX DC as a the LSC trigger but this did not work as there was no significant change in it when the arm was flashing. Eric then tried misaligning the ITM and using AS110 as a trigger - this worked. We then recompiled the ASS model to take AS110 as an input, and ran the dither alignment. After doing so, I measured the beam height at two points on the new endtable.

Bottom line:

  • The beam is roughly level across the table (along the North-South direction, within the precision to which I could place the irides and measure the height). The table has also been levelled pretty well...
  • The beam height is ~4.7" across the endtable

So the beam is about 0.7" higher relative to the endtable than we'd like it to be. What do we do about this?

  • Is it even possible to raise the table by 0.7" so we can have a level beam everywhere? Are there some constraints related to how the enclosure is attached to the window?
  • Are we okay with tolerating a solution where we keep the beam level at 4", and use Y10 and Y11 (see layout in elog 12060) to raise the beam by 0.7", and then have slightly higher posts for the optics downstream of this point?

I've also placed two irides extending the cavity axis on the endtable. These should be helpful in aligning the green to the arm eventually.

The new TMC 4' x 3' x4" optical table and enclosure is installed - aligned- leveled.

Atm2,  Picture is taken ~42" from the window at 3.75 camera height. The leveled table height is wthin 1/4 at the center of the window.

I think this is close enough to move on with the installation of the optics.

We can raise the loaded table in the future if it is needed.

Atm4, Optical table height to floor 33" at the south west corner

Atm3, Enclosure top cover transmission at 1064 nm, 1mm beam size, power level 157 mW, 0 degree incident angle,   T 1.3% Metal shield is required above 100 mW hitting the wall of the enclosure!

Atm5, window to enclosure Kapton seal

Attachment 1: ETMXbrs.png
ETMXbrs.png
Attachment 2: ETMX.jpg
ETMX.jpg
Attachment 3: IMG_0618.JPG
IMG_0618.JPG
Attachment 4: ETMX.jpg
ETMX.jpg
Attachment 5: ETMXwc.jpg
ETMXwc.jpg
  12066   Thu Apr 7 12:51:24 2016 gautamUpdateendtable upgradeBeam height differences

Steve has finished installing the enclosure on the new endtable. So Eric and I decided to try and lock the X arm and measure the beam height of the transmitted IR beam relative to the endtable. We initially thought of using POX DC as a the LSC trigger but this did not work as there was no significant change in it when the arm was flashing. Eric then tried misaligning the ITM and using AS110 as a trigger - this worked. We then recompiled the ASS model to take AS110 as an input, and ran the dither alignment. After doing so, I measured the beam height at two points on the new endtable.

Bottom line:

  • The beam is roughly level across the table (along the North-South direction, within the precision to which I could place the irides and measure the height). The table has also been levelled pretty well...
  • The beam height is ~4.7" across the endtable

So the beam is about 0.7" higher relative to the endtable than we'd like it to be. What do we do about this?

  • Is it even possible to raise the table by 0.7" so we can have a level beam everywhere? Are there some constraints related to how the enclosure is attached to the window?
  • Are we okay with tolerating a solution where we keep the beam level at 4", and use Y10 and Y11 (see layout in elog 12060) to raise the beam by 0.7", and then have slightly higher posts for the optics downstream of this point?

I've also placed two irides extending the cavity axis on the endtable. These should be helpful in aligning the green to the arm eventually.

  12065   Wed Apr 6 17:52:21 2016 gautamUpdateendtable upgradeFirst contact cleaning commenced

I've begun cleaning the optics that will eventually go back onto the newly installed X-endtable. We decided that First Contact was the way to go (as opposed to methanol drag wiping). Koji demonstrated the application of the (red) First Contact solution onto a 2" mirror - I then proceeded to work on the rest of the optics. We are broadly following the procedure in E1000079 - first one coat of First Contact solution is applied, then a small piece of PEEK is embedded by applying a second layer of solution over it (this will enable us to pull off the First Contact once we are ready - the plan is to do this after roughly placing the optic on the table. As of now, I've finished coating most of the optics that are part of the IR Transmon path - I will continue later in the evening.

The new endtable is almost ready for re-population. Steve just needs to shim the enclosure which will be done tomorrow morning. The game-plan as discussed at the meeting today is to first try and set up the IR Transmon path. This will allow us to verify that the endtable height is such that we can maintain a beam height of 4" everywhere on the table (I suspect we may have to compromise at some poing and do some fine adjustment of 1/4 to 1/2" somewhere though). It will also allow me to define the cavity axis relative to the table, which will be useful to place the green steering optics eventually. Doing this will be challenging though as right now, I can't see any of the arm flashes on the endtable using an IR card. Ideally, we want to somehow lock the X arm and then do the checks mentioned at the endtable, before beginning to put the endtable back together. 

  12064   Tue Apr 5 14:16:34 2016 gautamUpdateCDSBLRMS for optics suspensions - library block UPDATED

As discussed in a Wednesday meeting some time ago, we don't need to be writing channels from BLRMS filter modules to frames at 16k (we suspect this is leading to the frequent daqd crashes which were seen the last time we tried setting BLRMS up for all the suspensions). EricQ pointed out to me that there conveniently exists a library block that is much better suited to our purposes, called BLRMS_2k. I've replaced all the BLRMS library blocks in the sus_single_BLRMS library block that I made with there BLRMS_2k blocks. I need to check that the filters used by the BLRMS_2k block (which reside in /opt/rtcds/userapps/release/cds/common/src/BLRMSFILTER.c) are appropriate, after which we can give setting up BLRMS for all the suspensions a second try...

  12063   Tue Apr 5 11:42:17 2016 gaericqutamUpdateendtable upgradeTABLE REMOVAL

There is currently no table at the X end!

We have moved the vast majority of the optics to a temporary storage breadbord, and moved the end table itself to the workbench at the end. 

Steve says Transportation is coming at 1PM to put the new table in.

  12062   Tue Apr 5 08:55:51 2016 SteveUpdateSUSlocal EQ 3.1m

Local earth quake 3.1 magnitude in Valencia, Ca did not trip our suspensions.

 

Attachment 1: eq3.1Valencia.png
eq3.1Valencia.png
  12061   Mon Apr 4 15:04:14 2016 gautamUpdateendtable upgradeCOMPONENT REMOVAL


I'm planning to start removing components from the X endtable tomorrow morning at ~10AM - if anyone thinks I should hold off and do some further checks/planning, let me know before this so that I can do the needful.

  12060   Mon Apr 4 10:59:12 2016 gautamUpdateendtable upgradeproposed layout v3

I realized I had overlooked an important constraint in the layout, which is that the enclosure will have two supports that occupy some region of the table - these are denoted in blue in v3 of the layout (Attachment #1). I measured the dimensions for these from the existing Y-endtable. The main subsystem this has affected is the IR transmission monitors, but I've been able to move the photodiodes a little to accommodate this constraint.

I've also done the mode-matching calculations explicitly for the proposed new layout (Attachments #2 and #3, code in Attachment #4). While the layout was largely adopted from what Andres posted in this elog, I found that some of the parameters he used in his a la mode code were probably incorrect (e.g. distance between the 750mm lens and the ETM). More critically, I think the Gouy phase for the optimized solution in the same elog is more like 60 degrees. I found that I could get a (calculated) Gouy phase difference between the two PZT mirrors of ~81 degrees by changing the green path slightly, and making the two PZT mirrors Y7 and Y8 (instead of Y7 and Y11, for which the Gouy phase difference is more like 50 degrees). But this way the two steering mirrors are much closer to each other than they were before. Other misc. remarks about the mode matching calculations:

  • The beam diameter at the locations where the Faraday isolators should go is well below 5mm, the aperture size of the Faraday isolators
  • The calculated mode-matching efficiencies suggest that we don't need any cylindrical lenses though the mode from the NPRO is elliptical
  • Attachment #5 is a CAD drawing of the layout with all dimensions used for the mode-matching calculations included (although they are in inches)

These changes also necessitated minor changes to the transmitted IR beampath and the Oplev system, but these changes are minor. I've also switched the positions of the AUX IR power monitoring PD and the fiber coupler as suggested by Koji. The shutter has also been included. 

Attachment 1: ETMX_proposed_layout_v3.pdf
ETMX_proposed_layout_v3.pdf
Attachment 2: IR_modematch.pdf
IR_modematch.pdf
Attachment 3: Green_modematch.pdf
Green_modematch.pdf
Attachment 4: XendALaMode.zip
Attachment 5: ETMX_proposed_layout_dimensions.dwg
  12059   Fri Apr 1 13:11:26 2016 ericqUpdateWienerFilteringBilinear Noise Testing

I've been banging my head against bilinear noise subtraction, and figured I needed to test things on some real hardware to see if what I'm doing makes sense.


I ran the ASS dither alignment on the Y arm, which ensures that the beam spots are centered on both mirrors. 

I then drove ITMY in yaw with some noise bandpassed from 30-40 Hz. It showed the expected bilinear upconversion that you expect from angular noise on a centered beam, which you can see from 60-80 Hz below

I looked at the length signal, as the noise subtraction target, and the ITMY oplev yaw signal plus the transmon QPD yaw signal as witnesses.

There is some linear coupling to length, which means the the centering isn't perfect, and the drive is maybe large enough to displace it off center. However, the important part is the upconverted noise which is present only in the length signal. The QPD and oplev signals show no increased noise from 60-80Hz above the reference traces where no drive is applied

I then compared the multicoherence of those two angular witnesses vs. the multicoherence of the two (linear) witnesses plus their (bilinear) product. Including the bilinear term clearly shows coherence, and thereby subtraction potential, at the upconverted noise hump. 

So, it looks like the way I'm generating the bilinear signals and calculating coherence in my code isn't totally crazy.

Attachment 1: bilinear_drive.pdf
bilinear_drive.pdf
Attachment 2: 40m_bilin.pdf
40m_bilin.pdf
  12058   Thu Mar 31 19:49:31 2016 gautamUpdateendtable upgradeproposed layout v2

The major changes from the previous layout:

  1. I've depicted the Green reflected beam path more accurately - I approximately measured the angle of the rejected beam from the faraday from the Y-end setup. This looks like a workable solution, and is similar to what we have currently at the Y-end
  2. I've added some optics to monitor the DC power and RIN of the AUX laser
  3. I've added two lenses to the input path of the Oplev beam (the path is such that I think we can use the same lenses that are currently being used. 
  4. I've now drawn the beams in CAD so that is marginally neater.

To do:

  1. Post mode matching solutions for AUX laser to doubler and green beam to arm for this proposed layout (should be identical to what we have now, which at least according to the calculation is a good solution, but I will double check - I also need to quantify what the effect of the elliptical beam is)
  2. Check the Gouy phase of the transmitted IR beam at the QPD - we may need to change some lenses in this path. But I think the path as such is close enough (distance-wise) to what we have currently at the X end (after accounting for the fact that the new endtable edge will be closer to the ETM) so I don't expect this to be a show-stopper.

Does any part of this layout need a radical redesign? 

Attachment 1: ETMX_proposed_layout_v2.pdf
ETMX_proposed_layout_v2.pdf
  12057   Thu Mar 31 09:38:41 2016 SteveUpdateendtable upgradeY end 4x3 existing layout

Beam colors: 1064 nm red, 514 nm green and 633 nm yellow.

There should be room for lens in front of the pd at red3 and a mirror for alignment in the new layout.

This picture may help you how to improve the new ETMX 4' x 3' optical layout.

 

Attachment 1: ETMY4x3layout.jpg
ETMY4x3layout.jpg
  12056   Wed Mar 30 17:38:52 2016 gautamUpdateendtable upgradeX end table proposed layout

Attachment 1: This is a photo of the current X end table optical layout with the beampaths of the various sub-systems overlaid. For the labels, see Attachment #2.

Attachment 2: This is a summary of all the optical components that are currently being used. I've noted some things we may want to change when we effect the swap. The important ones are:

  • Switch out all 1" and 2" optic mounts which are not of the Polaris type to the Polaris type. I have checked that we have sufficient numbers of these in hand.
  • Adjust the collimating lens of the fiber collimating telescope to get a better mode
  • Many of the labels are probably outdated, now would be a good time to update them
  • For the mode-matching of the AUX IR into the doubling crystal, a la mode suggests a better (i.e. less sensitive to lens position) solution is effected with L2 as a 100mm fl lens rather than 88.3mm. I did not change this during the laser swap in order to minimize the number of components changed. Since we are doing a wholesale change now, it may not be a bad idea to swap this out as well. I have checked that we have a suitable AR1064 coated lens.
  • Some optics probably need to be cleaned...
  • PZT mirror 2 has a new mount ready that is the "correct" height so we don't have to keep using makeshift stacked posts.
  • The plan as it stands is to use the green coloured mount for the IR faraday (IO-5-1064-HP).

Have I missed anything important?

Attachment #3: I've made a CAD drawing of the proposed new layout and have overlaid the beampath in an amateur way because I couldn't figure OptoCad out - I figure this will suffice for now. I have adopted elements from the current Y-end layout, but have used Anders' mode-matching solution (same lenses, same positions of optics) to make sure we have good Guoy phase separation between the two PZT steering mirrors. Some notes:

  • I've tried to palce the optics for the AUX IR into the doubler and subsequent steering of green into the arm cavity as per the mode matching solution. These should be pretty accurate, and the layout suggests we have some room to maneuver 
  • The Green REFL beampath is exaggerated but I think we have enough room to place Y16 appropriately and steer the reflected beam into the PDA36A
  • We need two more 1" 1064nm coated mirrors for the initial steering into the doubling oven, I have checked we have these in hand.
  • The IR pickoff into the fiber coupler may change somewhat once we change the mode and redo the mode-matching calculations. But again, I think we have sufficient room to implement a workable solution.
  • After accounting for the fact that the new endtable will be a little closer to the vacuum chamber, Y12 in the proposed layout will be ~10cm further away from ETMX than it is currently. But as discussed at the meeting today, the Rayleigh range of the green beam should be large enough here such that this shouldn't be a significant change.

Steve says the table is ready - so if we are happy with this layout, we can move forward...

Attachment 1: ETMX_3x2.JPG
ETMX_3x2.JPG
Attachment 2: layout_details_20160328.pdf
layout_details_20160328.pdf
Attachment 3: ETMX_proposed_layout.pdf
ETMX_proposed_layout.pdf
  12055   Wed Mar 30 16:40:24 2016 ericqUpdateLSC2016 vs 2010

I haven't found any data files for the DARM spectrum of the previous generation of 40m, but with some GIMP-fu, I have plotted Monday's spectrum (green) on top of one of the figures from Rob's thesis.

  12054   Wed Mar 30 11:35:24 2016 steveUpdatesafetyBS visitor's viewport is protected with lexan

Quote:
The four horizontal viewports of arms are protected
by 3/8" thick, 8.5" OD Lexan disk of MR10 Polycarbonate.

ITMX, ETMX, ITMY and ETMY ccd cameras are not focused now.


BS visitor's viewport glass is now covered with Lexan MR10

Note:
this Lexan cover is in vertical orientation so becomes lose when the black anodized cover is removed.
It needs to be held in place

while it's housing is taken off.
  12053   Tue Mar 29 03:16:21 2016 ericqUpdateLSCDRFPMI Locked Once More

[ericq, Gautam]

Three RF-only locks longer than a minute tonight, out of 5 total attempts. 

Last week, I determined that the beam spot on the RF POP PD is too large. This still needs to be fixed. I updated the ASS model to use REFLDC as a PRCL dither error signal; it works. 

There seems to be some excess angular motion of ETMY tonight. This is evident in the oplev spectra (as compared to ETMX), and the GTRY camera, and even the retroreflected beam from a misalgined ETMY on the ITMY face when the PRC is carrier locked.

Gautam and I mostly focused on setting up the CAL-DARM_CINV block to produce this (mostly) calibrated spectrum starting from GPS 1143274087. [Darm on unwhitened AS55, DRMI on 3F, one CARM boost]

Here are the control and error signal spectra:

[DTT files attached]

Note to self: archive some of this data 

Attachment 1: 2016-03-29_calibdarm.pdf
2016-03-29_calibdarm.pdf
Attachment 2: 2016-03-29_DRFPMI_errctrl.pdf
2016-03-29_DRFPMI_errctrl.pdf
Attachment 3: DRFPMI_DTT.zip
  12052   Mon Mar 28 22:16:44 2016 KojiUpdateGeneralNew WiFi router

I configured three more mini wifi extender. They are ready to use.

We should add these to the host table (I forgot where it is)

192.168.113.233 NETGEAR_EX3700_1
192.168.113.234 NETGEAR_EX3700_2
192.168.113.235 NETGEAR_EX3700_3
192.168.113.236 NETGEAR_EX3700_4

  12051   Mon Mar 28 10:43:18 2016 SteveUpdateendtable upgradeETMX 4'x2' optical table pictures

 

 

Attachment 1: ETMX4x2Layout.jpg
ETMX4x2Layout.jpg
Attachment 2: ETMX_4x2.JPG
ETMX_4x2.JPG
Attachment 3: ETMXopt_4x2.JPG
ETMXopt_4x2.JPG
  12050   Mon Mar 28 08:30:09 2016 SteveUpdateSUSPRM damping restored

Recent  EQ 4.8 mag San Felipe, Mexico trips PRM sus damping.

PRM damping restored. PMC locked.

Attachment 1: shaking.png
shaking.png
  12049   Sat Mar 26 18:28:24 2016 KojiUpdateelogelogd flakiness

Elogd have been restarted several times today because it died everytime I submit something.
Here is the copy of the log.


GET /OMC_Lab/255?cmd=loc&value=Submit HTTP/1.1
Returned 6 bytes
GET /40m/elog.rdf HTTP/1.1
Returned 17109 bytes
TCP connection #1 on socket 5 closed
POST /OMC_Lab/ HTTP/1.1
Returned 20 bytes
GET /OMC_Lab/255 HTTP/1.1
Returned 53721 bytes
GET /ckeditor/skins/moono/images/arrow.png HTTP/1.1
Returned 489 bytes
POST /OMC_Lab/ HTTP/1.1
*** buffer overflow detected ***: /export/home/elog/elog/elogd terminated
======= Backtrace: =========
/lib/x86_64-linux-gnu/libc.so.6(__fortify_fail+0x37)[0x7f1435639e57]
/lib/x86_64-linux-gnu/libc.so.6(+0x108d50)[0x7f1435638d50]
/lib/x86_64-linux-gnu/libc.so.6(+0x1081b9)[0x7f14356381b9]
/lib/x86_64-linux-gnu/libc.so.6(_IO_default_xsputn+0xdd)[0x7f14355ab0cd]
/lib/x86_64-linux-gnu/libc.so.6(_IO_vfprintf+0x25a8)[0x7f143557ac18]
/lib/x86_64-linux-gnu/libc.so.6(__vsprintf_chk+0x94)[0x7f1435638254]
/lib/x86_64-linux-gnu/libc.so.6(__sprintf_chk+0x7d)[0x7f143563819d]
/export/home/elog/elog/elogd[0x426405]
/export/home/elog/elog/elogd[0x473b7f]
/export/home/elog/elog/elogd[0x4abfb2]
/export/home/elog/elog/elogd[0x4ad7fb]
/export/home/elog/elog/elogd[0x4b0af5]
/export/home/elog/elog/elogd[0x4b1eb9]
/export/home/elog/elog/elogd[0x403568]
/lib/x86_64-linux-gnu/libc.so.6(__libc_start_main+0xed)[0x7f143555176d]
/export/home/elog/elog/elogd[0x404299]
======= Memory map: ========
00400000-004e6000 r-xp 00000000 fc:00 10361276                           /export/home/elog/elog-3.0.d/elogd
006e5000-006e6000 r--p 000e5000 fc:00 10361276                           /export/home/elog/elog-3.0.d/elogd
006e6000-007c6000 rw-p 000e6000 fc:00 10361276                           /export/home/elog/elog-3.0.d/elogd
007c6000-0173d000 rw-p 00000000 00:00 0
0214d000-02656000 rw-p 00000000 00:00 0                                  [heap]
7f14342f8000-7f143430d000 r-xp 00000000 fc:00 2883628                    /lib/x86_64-linux-gnu/libgcc_s.so.1
7f143430d000-7f143450c000 ---p 00015000 fc:00 2883628                    /lib/x86_64-linux-gnu/libgcc_s.so.1
7f143450c000-7f143450d000 r--p 00014000 fc:00 2883628                    /lib/x86_64-linux-gnu/libgcc_s.so.1
7f143450d000-7f143450e000 rw-p 00015000 fc:00 2883628                    /lib/x86_64-linux-gnu/libgcc_s.so.1
7f143450e000-7f14348cd000 rw-p 00000000 00:00 0
7f1434a34000-7f1434d39000 r--p 00000000 fc:00 530477                     /usr/lib/locale/locale-archive
7f1434d39000-7f1434d4f000 r-xp 00000000 fc:00 655527                     /usr/local/lib/libz.so.1.2.8
7f1434d4f000-7f1434f4e000 ---p 00016000 fc:00 655527                     /usr/local/lib/libz.so.1.2.8
7f1434f4e000-7f1434f4f000 r--p 00015000 fc:00 655527                     /usr/local/lib/libz.so.1.2.8
7f1434f4f000-7f1434f50000 rw-p 00016000 fc:00 655527                     /usr/local/lib/libz.so.1.2.8
7f1434f50000-7f1434f52000 r-xp 00000000 fc:00 2883655                    /lib/x86_64-linux-gnu/libdl-2.15.so
7f1434f52000-7f1435152000 ---p 00002000 fc:00 2883655                    /lib/x86_64-linux-gnu/libdl-2.15.so
7f1435152000-7f1435153000 r--p 00002000 fc:00 2883655                    /lib/x86_64-linux-gnu/libdl-2.15.so
7f1435153000-7f1435154000 rw-p 00003000 fc:00 2883655                    /lib/x86_64-linux-gnu/libdl-2.15.so
7f1435154000-7f1435307000 r-xp 00000000 fc:00 2883609                    /lib/x86_64-linux-gnu/libcrypto.so.1.0.0
7f1435307000-7f1435506000 ---p 001b3000 fc:00 2883609                    /lib/x86_64-linux-gnu/libcrypto.so.1.0.0
7f1435506000-7f1435521000 r--p 001b2000 fc:00 2883609                    /lib/x86_64-linux-gnu/libcrypto.so.1.0.0
7f1435521000-7f143552c000 rw-p 001cd000 fc:00 2883609                    /lib/x86_64-linux-gnu/libcrypto.so.1.0.0
7f143552c000-7f1435530000 rw-p 00000000 00:00 0
7f1435530000-7f14356e4000 r-xp 00000000 fc:00 2884139                    /lib/x86_64-linux-gnu/libc-2.15.so
7f14356e4000-7f14358e3000 ---p 001b4000 fc:00 2884139                    /lib/x86_64-linux-gnu/libc-2.15.so
7f14358e3000-7f14358e7000 r--p 001b3000 fc:00 2884139                    /lib/x86_64-linux-gnu/libc-2.15.so
7f14358e7000-7f14358e9000 rw-p 001b7000 fc:00 2884139                    /lib/x86_64-linux-gnu/libc-2.15.so
7f14358e9000-7f14358ee000 rw-p 00000000 00:00 0
7f14358ee000-7f1435943000 r-xp 00000000 fc:00 2884155                    /lib/x86_64-linux-gnu/libssl.so.1.0.0
7f1435943000-7f1435b42000 ---p 00055000 fc:00 2884155                    /lib/x86_64-linux-gnu/libssl.so.1.0.0
7f1435b42000-7f1435b45000 r--p 00054000 fc:00 2884155                    /lib/x86_64-linux-gnu/libssl.so.1.0.0
7f1435b45000-7f1435b4c000 rw-p 00057000 fc:00 2884155                    /lib/x86_64-linux-gnu/libssl.so.1.0.0
7f1435b4c000-7f1435b6e000 r-xp 00000000 fc:00 2884145                    /lib/x86_64-linux-gnu/ld-2.15.so
7f1435d57000-7f1435d5c000 rw-p 00000000 00:00 0
7f1435d6a000-7f1435d6e000 rw-p 00000000 00:00 0
7f1435d6e000-7f1435d6f000 r--p 00022000 fc:00 2884145                    /lib/x86_64-linux-gnu/ld-2.15.so
7f1435d6f000-7f1435d71000 rw-p 00023000 fc:00 2884145                    /lib/x86_64-linux-gnu/ld-2.15.so
7ffd85795000-7ffd85997000 rw-p 00000000 00:00 0                          [stack]
7ffd859b2000-7ffd859b4000 r-xp 00000000 00:00 0                          [vdso]
ffffffffff600000-ffffffffff601000 r-xp 00000000 00:00 0                  [vsyscall]
er_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
Received unknown cookie "__utma"
Received unknown cookie "_ga"
Received unknown cookie "__unam"
Received unknown cookie "__utma"
Received unknown cookie "tk_ni"
Received unknown cookie "ajs_user_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
Received unknown cookie "cache_disable"
Received unknown cookie "NO_CACHE"
Received unknown cookie "__utma"
Received unknown cookie "_ga"
Received unknown cookie "__unam"
Received unknown cookie "__utma"
Received unknown cookie "tk_ni"
Received unknown cookie "ajs_user_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
Received unknown cookie "__utma"
Received unknown cookie "_ga"
Received unknown cookie "__unam"
Received unknown cookie "__utma"
Received unknown cookie "tk_ni"
Received unknown cookie "ajs_user_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
Received unknown cookie "__utma"
Received unknown cookie "_ga"
Received unknown cookie "__unam"
Received unknown cookie "__utma"
Received unknown cookie "tk_ni"
Received unknown cookie "ajs_user_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
Received unknown cookie "__utma"
Received unknown cookie "_ga"
Received unknown cookie "__unam"
Received unknown cookie "__utma"
Received unknown cookie "tk_ni"
Received unknown cookie "ajs_user_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
  12048   Fri Mar 25 23:54:04 2016 ranaUpdatePEMGuralp Seismometers

Something seems not right. The Guralp response should be flat in velocity from 0.05-30 Hz. Why is there any feature at 1 Hz? Saturation of some kind?

  12047   Fri Mar 25 19:17:28 2016 NikhilUpdatePEMGuralp Seismometers

Calibration of Guralp Seismometers

Objective

  • Estimate transfer functions of Guralp A ( near ETMX) and Guralp B ( near ETMY)
  • Calibrate the instruments by estimating Velocity Sensitity Parameter
  • Convert previously measured Voltage Spectrum to Velocity Spectrum

Instruments Used

  • Guralp CMG-40 T Seimometers  : Guralp A (Serial Number: T4Q17)
  • Guralp CMG-40 T Seimometers  : Guralp B (Serial Number: T4157)
  • Guralp Handheld Control Unit (HCU)
  • FFT Spectrum Analyzer: Model SR785: 2 Channel Dynamic Signal Analyzer
  • Oscilloscope: TDS 3014B
  • Function Generator: DS 345

Procedure & Results

Sinusoidal current of known frequency and amplitude was injected to the Seismometer calibration coil using signal generator and handheld control unit & corresponding Magnitude and Phase response were recorded.  For  Guralp B, system response was also estimated with a FFT Spectrum Analyzer. 

 

    

Frequnecy Range: 0.1 Hz to 45 Hz.

Equivalent Input Velocity was derived from the Input Voltage measurements using the relation: v = V/ (2*pi*f*R*K) , V is the peak to peak Calibration Signal voltage, f is the calibration signal frequency, R is the calibration resistor and K is the feedback coil constant.  [See Appendix for R & K values]                                     

Velocity Sensitity at the required frequency is obtained by dividing the Output Response Voltage by the Equivalent Input Velocity.

               

The obtained Velocity Sensitivity is used to convert the recorded Volatge PSD to Velocity PSD as shown below. The obtained results are compared to gloabl high noise model [NHNM] and USGS New Low Noise Model [NLNM,Peterson 1993] which gives the lowest observed vertical seismic noise levels across the seismic frequency band. Plot legend NLNM shows both the high & low levels.

 

                                                                    Guralp A [X Arm] Low Velocity Output                                          

                  

                                                                    Guralp B [Y Arm] Low Velocity Output                                          

            

                                                                            DTT Power Spectrum                                                             

Both the Seismometers were connected to the 40 M Control and Data Acquisition System (CDS) and Power Spectrum was estimated for the Vertical, North/South & East/West Channels using Diagnostic Test Tool (DTT) software.

                                                            

Comments

  •  The transfer function from Guralp A [ETMX] looks similar to that of Guralp B [ETMY] in both magnitude and phase but with a lower gain.                                                                                                                                                                                                  
  • Velocity Sensitivity of Guralp A is comparable to the value provided in the Calibration Data Sheet [~ 400] for all the channels [Vertical, North/South, East/West] after 1 Hz. For Guralp B, Velocity Sensitivity is a factor of 2.5 higher [all channels] than the specification [~ 400] after 1 HZ.Below 1 Hz Sensitivity drops down for both sensors. I am not ruling out a missing common factor in the calculation, but anyway, test shows that Guralp B has ~2.5 times better Velocity Sensitivity than Guralp A.                                                                                                                                                               
  • The Calibrated Seismic Velocity Spectrum for Guralp B is within the Globally Observed High and Low Noise Seismic Spectrum while Guralp A's Spectrum is more noisier above 1 Hz [Anthropogenic Activity normally contributes the most in 1 Hz to 10 Hz frequency band].                                                                                                                                                                                                                                                                          
  • Concurrently acquired Power Spectrum using DTT [Diagnostic Test Tools] shows that Guralp A Spectrum behaves rather strangely. The system response seems to be completely different from the one we obtained locally using signal generator. While Guralp B functionality seems normal. One reason for this erratic beahvior might be faulty cables used for data acquisition from Guralp A. This needs to be verified.                                                                                                                        

Appendix

                                                                            CMG-40T Guralp A Calibration Sheet                                                           

  Velocity Output: V/m/s (Differential) Mass Position Output (Acceleration Output) : V/m/s^2  Feedback Coil Constant : Amp/m/s^2
Vertical 2 x 400 19 0.00397
North/South 2 x 398 23 0.00486
East/West 2 x 401 23 0.00479                 

Calibration Resistor: 51000

                                                                            CMG-40T Guralp B Calibration Sheet                                                           

  Velocity Output: V/m/s (Differential) Mass Position Output (Acceleration Output) : V/m/s^2  Feedback Coil Constant : Amp/m/s^2
Vertical 2 x 401 19 0.00408
North/South 2 x 400 20 0.00421
East/West 2 x 404 22

0.00466

Calibration Resistor: 51000 

Attachment 3: VelSens_XArm_Guralp_A.png
VelSens_XArm_Guralp_A.png
Attachment 4: VelSens_XArm_Guralp_A.png
VelSens_XArm_Guralp_A.png
Attachment 5: VelSens_YArm_Guralp_B.png
VelSens_YArm_Guralp_B.png
Attachment 8: Vel_PSD_XArm_GurB_E.png
Vel_PSD_XArm_GurB_E.png
Attachment 9: Vel_PSD_XArm_GurB_N.png
Vel_PSD_XArm_GurB_N.png
Attachment 16: Guralp_PowerSpectrum.pdf
Guralp_PowerSpectrum.pdf
  12046   Thu Mar 24 08:20:52 2016 SteveUpdatePEMGuralp-A calibration sheet

Calibration Data

All Guralp instruments and digitisers are provided with calibration documentation. Should you require a copy of calibration information for any product, email caldoc@guralp.com with the serial number of the product in the subject field and calibration information will be sent to you through email.

See data in the 40m wiki

 

  12045   Thu Mar 24 07:56:09 2016 SteveUpdateCalibration-RepairNO Noise Eater for 1W Innolight

1W Innolight is NOT getting Noise Eater as it was decided yesterday at the 40m meeting. Corrected 3-25-2016

Repair quote with adding noise eater is in 40m wiki

Quote:

 

Quote:
Quote:

After adjusting the alignment of the two beams onto the PD, I managed to recover a stronger beatnote of ~ -10dBm. I managed to take some measurements with the PLL locked, and will put up a more detailed post later in the evening. I turned the IMC autolocker off, turned the 11MHz Marconi output off, and closed the PSL shutter for the duration of my work, but have reverted these to their nominal state now. The are a few extra cables running from the PSL table to the area near the IOO rack where I was doing the measurements from, I've left these as is for now in case I need to take some more data later in the evening...I

Innolight 1W 1064nm, sn 1634 was purchased in 9-18-2006 at CIT. It came to the 40m around 2010

It's diodes should be replaced, based on it's age and performance.

RIN and noise eater bad. I will get a quote on this job.

The Innolight Manual frequency noise plot is the same as Lightwave' elog 11956

Diagnoses from Glasglow:

“So far we have analyzed the laser. The pump diode is degraded. Next we would replace it with a new diode. We would realign the diode output beam into the laser crystal. We check all the relevant laser parameters over the whole tuning range. Parameters include single direction operation of the ring resonator, single frequency operation, beam profile and others. If one of them is out of spec, then we would take actions accordingly. We would also monitor the output power stability over one night. Then we repackage and ship the laser.”

 

ELOG V3.1.3-